# Entanglement dynamics in intensity-dependent double Jaynes-Cummings model for squeezed coherent thermal states

### v6

@@ Abstract
In this paper, the entanglement dynamics of different subsystems such as atom-atom, atom-field and field-field with radiation field in squeezed coherent thermal states for the intensity-dependent double Jaynes-Cummings model (IDDJCM) and double Jaynes-Cummings model (DJCM) are investigated. The effects of squeezed and thermal photons on entanglement are examined, revealing their complementary roles in shaping the entanglement behavior in both models. One of the main features of the double Jaynes-Cummings model is the observation of entanglement sudden death for every subsystem. The effects of various interactions such as Ising interaction, single photon exchange interaction and dipole-dipole interaction on entanglement dynamics are studied. The effects of detuning, Kerr-nonlinearity on the entanglement dynamics are investigated for every subsystem. It is noticed that proper choice of the interactions parameters, detuning and Kerr-nonlinearity effectively removes entanglement deaths from the dynamics.
@@ 1 Introduction
It is well known that the Jaynes-Cummings (J-C) model is one of the most successful models to describe the atom-field interaction and to study the dynamics of various quantum optical aspects. Entanglement [1] is one of the important quantities which can be investigated using the Jaynes-Cummings model. In recent years, entanglement has become ubiquitous and necessary for quantum applications. It has become a resource for quantum teleportation [2], super dense coding [3], entanglement swapping, cryptography etc. In many quantum optical systems such as in ion traps [4, 5, 6, 7], cavity quantum electrodynamics (CQED) [8, 9, 10, 11, 12, 13, 14, 15], circuit quantum electrodynamics [16, 17] etc., the entanglement is studied. All these physical systems are a mixture of atoms and fields. Jaynes-Cummings model provides a platform to study entanglements in these systems.
One of the successful extensions of the usual Jaynes-Cummings model is the double Jaynes-Cummings model (DJCM), which was first introduced by Yonac et. al., [18, 19, 20, 21, 22, 23, 24, 25, 26]. This model can be used to study the properties of entanglement in a more extended quantum optical systems which are mentioned before. A striking feature of the DJCM is the sudden disappearance or death of entanglement for a certain duration. This phenomenon of disappearance of entanglement is termed as entanglement sudden death (ESD). After Yonac’s work, there are studies on various systems with this model [27, 28, 29, 30]. In particular in Ref. [29], the authors have studied entanglement dynamics of the coherent and squeezed vacuum states and the atoms. They have shown that the factors such as atomic spontaneous decay rate, cavity decay rate and detuning have significant effects on the entanglement sudden death. To study the effects of interaction, the authors in Ref. [28, 31] considered cavities with Ising type and photon exchange interactions between the cavities. The notable features of entanglement decay, sudden rebirth and sudden death have been investigated in the presence of intrinsic decoherence in Ref. [32]. In another work, Laha [33] has considered a double Jaynes-Cummings model to investigate the role of a beam-splitter, dipole-dipole interaction as well as Ising type interactions. The photonic modes in this work are considered to be either the photonic vacuum state, or the coherent states or the thermal states and the entanglements investigated are the qubit-qubit and the oscillator-oscillator type.
@@
After Yonac’s work on double Jaynes-Cummings model and the entanglement sudden death, various systems with double Jaynes-Cummings model have been studied; however, there are very few studies on the intensity-dependent double Jaynes-Cummings model (IDDJCM). In [43] Xie Qin and Fang Mao-Fa have investigated the entanglement dynamics of IDDJCM with two different initial radiation fields. One is a coherent state and another is a squeezed vacuum state. Motivated by these works, in this paper, we study the entanglement dynamics for both double Jaynes-Cummings and intensity-dependent double Jaynes-Cummings models with the radiation fields in squeezed coherent thermal states (SCTS), and, with the atoms in a Bell state. One of the main objectives of this work is to investigate the effects of squeezed photons and thermal photons on the entanglement in a coherent background for DJCM and IDDJCM. Studying the effects of different kinds of interactions on the entanglement dynamics in the system is also an objective. In this paper, the effects of diverse interactions such as single photon exchange interaction, Ising interaction and dipole-dipole interaction on entanglement are investigated. We also investigate the effects of detuning, Kerr-nonlinearity [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] etc. Earlier, in Ref.[56], the authors have studied the effects of the squeezed and thermal photons on entanglements in Jaynes-Cummings model. Here, it was shown how the tussling between “classical noise” (thermal photons) and “quantum noise” (squeezed photons) affects the quantum optical properties of the radiation field, atomic inversion and entanglement dynamics for Jaynes-Cummings model. The work presented here is a natural extension of our previous work to investigate the effects of the perpetual tussle between thermal and squeezed photons on the entanglement dynamics of various subsystems in DJCM and IDDJCM.
@@ 2.1 Photonic state
$\hat{\rho}_{\text{th}}=\frac{1}{1+\bar{n}_{th}}\sum_{n=0}^{\infty}\left(\frac{\bar{n}_{th}}{\bar{n}_{th}+1}\right)^{n}\ket{n}\bra{n},$ (4)
@@
×$\times\big{|}H_{q}((2Y)^{-1/2}\tilde{Z})\big{|}^{2},$ (7)
where $\pi Q(0)=R(0,0)$; $R$ is Glauber’s $R$-function[60];
where
$X$ $=\bar{n}_{th}+(2\bar{n}_{th}+1)(\sinh r)^{2},$ (9)
$Y$ $=-(2\bar{n}_{th}+1)e^{i\varphi}\sinh r\cosh r,$ (10)
[…]
and
$\tilde{X}$ $=\frac{X(1+X)-|Y|^{2}}{(1+X)^{2}-|Y|^{2}},$ (13)
$\tilde{Y}$ $=\frac{Y}{(1+X)^{2}-|Y|^{2}},$ (14)
$\tilde{Z}$ $=\frac{(1+X)Z+YZ^{*}}{(1+X)^{2}-|Y|^{2}}.$ (15)
$\tilde{X}$ $=\frac{\bar{n}_{th}(\bar{n}_{th}+1)}{\bar{n}_{th}^{2}+(\bar{n}_{th}+\frac{1}{2})[1+\cosh(2r)]}$ (16)
$\tilde{Y}$ $=-\frac{e^{i\varphi}(\bar{n}_{th}+\frac{1}{2})\sinh(2r)}{\bar{n}_{th}^{2}+(\bar{n}_{th}+\frac{1}{2})[1+\cosh(2r)]}$ (17)
$\tilde{Z}$ $=\frac{Z[\frac{1}{2}+(\bar{n}_{th}+\frac{1}{2})\cosh r]-Z^{*}e^{i\varphi}(\bar{n}_{th}+\frac{1}{2})\sinh 2r}{\bar{n}_{th}^{2}+(\bar{n}_{th}+\frac{1}{2})[1+\cosh(2r)]}$
and $H_{q}$ is the Hermite polynomial. It is defined as
@@ 2.2 Atomic state
In the double Jaynes-Cummings model, there are two two-level atoms, one in cavity a and the other in cavity b, respectively. These two-level systems are initially prepared in either as a pure state or as a mixed state. For the pure state a maximally entangled Bell state is considered. In this paper, we use the entangled state of the form
@@
where $|e_{\text{A}}\rangle$, $|e_{\text{B}}\rangle$ and $|g_{\text{A}}\rangle$,$|g_{\text{B}}\rangle$ correspond to the excited and ground states of the atom A and atom B respectively.
@@ 2.3 The Hamiltonian of the system
The initial state vector of the atom-field system is
[…]
The basis states for this atom-field system are |$\ket{e,g,n,m},\ket{g,e,n,m},\ket{g,g,n+1,m},\ket{g,g,n,m-1},$ |$\ket{e,e,n-1,m},\ket{e,e,n,m-1},\ket{g,e,n+1,m-1},\ket{e,g,n-1,m+1}$. To get $\psi(t)$ we need to solve the Schrodinger equation for IDDJCM and DJCM. For this purpose, let us solve the single Jaynes-Cummings models first. Let the initial state of the atom-field system in the single J-C model be
|$\ket{\psi_{I}(0)}=\ket{e}\otimes\ket{n}.$ (26)
After time evolution under the single Jaynes-Cummings Hamiltonian, the state becomes
|$\ket{\psi_{I}(t)}=x_{1}(t)\ket{e}\ket{n}+x_{2}(t)\ket{g}\ket{n+1},$ (27)
where $x_{1}(t)$ and $x_{2}(t)$ are the probability amplitudes for the system to be found in |$\ket{e}\ket{n}$ and |$\ket{g}\ket{n+1}$ states respectively. After solving the Schrödinger equation for this state with the initial conditions $x_{1}(0)=1$ and $x_{2}(0)=0$, we get
$x_{1}(t)=\cos({\lambda\sqrt{n+1}\,t}),$ (28)
$x_{2}(t)=-i\sin(\lambda\sqrt{n+1}\,t).$ (29)
For the intensity-dependent single J-C Hamiltonian
$x_{1}(t)=\cos({\lambda(n+1)\,t}),$ (30)
$x_{2}(t)=-i\sin(\lambda(n+1)\,t).$ (31)
If the system starts from
|$\ket{\psi_{I}(0)}=\ket{g}\otimes\ket{n},$ (32)
after evolution, the state becomes
[…]
As before, solving for the initial conditions $y_{1}(0)=1$ and $y_{2}(0)=0$, we get
$y_{1}(t)=\cos(\lambda\sqrt{n}\,t),$ (34)
@@
$y_{2}(t)=-i\sin(\lambda\sqrt{n}\,t).$ (35)
In the intensity-dependent case
$y_{1}(t)=\cos(\lambda n\,t),$ (36)
and
$y_{2}(t)=-i\sin(\lambda n\,t).$ (37)
In the double Jaynes-Cummings model, the initial state for the atom-field system is
[…]
where we consider the atomic states corresponding to the cavities a and b are entangled. So, after time evolution the initial state in Eq. (38) evolves to
|$\ket{\psi(t)}$ $=$ $\cos\theta\sum_{n}c_{n}(x_{1}(t)\ket{e,n}+x_{2}(t)\ket{g,n+1})\otimes\sum_{m}d_{m}(y_{1}(t)\ket{g,m}+y_{2}(t)\ket{e,m-1})$
$+\sin\theta\sum_{n}c_{n}(x_{1}(t)\ket{e,n}+x_{2}(t)\ket{g,n+1})\otimes\sum_{m}d_{m}(y_{1}(t)\ket{g,m}+y_{2}(t)\ket{e,m-1}),$
which can be rewritten as
|$\ket{\psi(t)}$ $=$ $\sum_{n,m=0}^{\infty}a_{1}(n,m,t)\ket{e,g,n,m}+a_{2}(n,m,t)\ket{e,e,n,m-1}+a_{3}(n,m,t)\ket{g,g,n+1,m}$ (40)
$+a_{4}(n,m,t)\ket{g,e,n+1,m-1}+a_{5}(n,m,t)\ket{g,e,n,m}+a_{6}(n,m,t)\ket{g,g,n,m+1}$
$+a_{7}(n,m,t)\ket{e,e,n-1,m}+a_{8}(n,m,t)\ket{e,g,n-1,m+1},$
where the factors $a_{i}$’s are as given below:
$a_{1}(n,m,t)$ $=$ $\cos\theta\sum_{n}c_{n}x_{1}(t)\sum_{m}d_{m}y_{1}(t);\qquad a_{2}(n,m,t)=\cos\theta\sum_{n}c_{n}x_{2}(t)\sum_{m}d_{m}y_{2}(t);$
$a_{3}(n,m,t)$ $=$ $\cos\theta\sum_{n}c_{n}x_{2}(t)\sum_{m}d_{m}y_{1}(t);\qquad a_{4}(n,m,t)=\cos\theta\sum_{n}c_{n}x_{2}(t)\sum_{m}d_{m}y_{2}(t);$
$a_{5}(n,m,t)$ $=$ $\sin\theta\sum_{n}c_{n}y_{1}(t)\sum_{m}d_{m}x_{1}(t);\qquad a_{6}(n,m,t)=\sin\theta\sum_{n}c_{n}y_{1}(t)\sum_{m}d_{m}x_{2}(t);$
$a_{7}(n,m,t)$ $=$ $\sin\theta\sum_{n}c_{n}y_{2}(t)\sum_{m}d_{m}x_{1}(t);\qquad a_{8}(n,m,t)=\sin\theta\sum_{n}c_{n}y_{2}(t)\sum_{m}d_{m}x_{2}(t).$
From the knowledge of the time evolved state, one can construct the time evolved density matrix
$\hat{\rho}(t)=\ket{\psi(t)}\bra{\psi(t)}$ | which can be used to calculate the entanglement between different subsystems like the atom-atom, the atom-field and the field-field subsystems.
@@ 2.4 Entanglement measures
where $\Lambda_{i}(i=1,2,3,4)$ are the decreasingly ordered square roots of the eigenvalues of the matrix $\hat{\rho}\left(\hat{\sigma}_{y}^{\text{A}}\otimes\hat{\sigma}_{y}^{\text{B}}\right)\hat{\rho}^{*}\left(\hat{\sigma}_{y}^{\text{A}}\otimes\hat{\sigma}_{y}^{\text{B}}\right)$ and $\hat{\rho}$ is the two qubit atom-atom reduced density matrix. The value of concurrence lies in the range $0\leq C\leq 1$, where $C=0$ implies a separable state and $C=1$ denotes a maximally entangled state. Though concurrence can be used to compute entanglement in both pure and mixed states, if the system is $2\otimes 2$ dimensions. So, for higher dimensional systems we need to use other measures. In particular, when we consider the atom-field subsystem we are looking at $2\otimes\infty$ and bipartite continuous variable systems. For these systems it is convenient to use the negativity [62] which is defined as
@@ 3 Entanglement dynamics in IDDJCM and DJCM
Figure 2: All the entanglements in Figure 1 are plotted in a single plot for various $n_{s}$.
@@
In this section, we study the entanglement dynamics of the atom-field system with radiation field in SCTS with atoms in a Bell state. Fig. 1 represents the entanglement dynamics for IDDJCM. To study the dynamics, $\bar{n}_{c}=2$, $\bar{n}_{th}=1.0$ and $\bar{n}_{s}=0.1,0.3,0.5,1.0$ have been taken. It is already shown in earlier studies of the intensity-dependent Jaynes-Cummings model that the dynamics of the system show a periodic nature [43]. From Fig. 1, it is observed that $C(t)$ for the atom-atom subsystem and $N(t)$ for all the other three subsystems show periodic dynamics. From Fig. 1(a), we notice that the dynamics of $C(t)$ does not change considerably with increasing squeezed photons. Lengths of the ESDs in the dynamics change very little with the increment in $\bar{n}_{s}$ and the heights of the peaks do not change at all. This is also exactly the same for the field-field subsystem in Fig. 1(d). For atom A-field a subsystem in Fig. 1(b), the amplitude of $N(t)$ increases keeping the shape of the dynamics almost similar; however, there is no change in ESDs. In Fig. 1(c), for atom A-field b subsystem, peak heights decrease and lengths of ESDs also decrease with increasing $\bar{n}_{s}$. So, for field-field subsystem, increase in number of squeezed photons does not affect the $C(t)$ and $N(t)$. The dynamics almost superimpose on one another. Fig. 2 shows how the entanglement is shared with respect to time between different subsystems for a fixed $\bar{n}_{s}$.
In Fig. 3, the entanglement dynamics for DJCM are plotted. The dynamics are very different from the dynamics obtained in IDDJCM. The periodic nature of entanglement (like periodic pulse) is not observed for DJCM. In the case of atom-atom entanglement, $C(t)$ starts from the highest value and then falls very sharply. $C(t)$ shows a large ESD and the duration of ESD increases with increasing value of $\bar{n}_{s}$. The effects of thermal photons are also visible here. In the case of squeezed coherent state [63, 64, 57, 58, 65, 59, 66, 67, 68, 69, 70, 71, 72, 73], smaller peaks are present in the dynamics of $C(t)$ and the height of the peaks $\lambda t=8$ were bigger. Those smaller peaks disappear and the heights of the aforementioned peaks also come down. In a sense, the thermal photons destroy the entanglement in atom-atom subsystem.

### v7

@@ Abstract
In this paper, the entanglement dynamics of different subsystems such as atom-atom, atom-field and field-field with radiation field in squeezed coherent thermal states for the intensity-dependent double Jaynes-Cummings model (IDDJCM) and double Jaynes-Cummings model (DJCM) are investigated. The effects of squeezed and thermal photons on entanglement are examined, revealing their complementary roles in shaping the entanglement behavior in both models. One of the main features of the double Jaynes-Cummings model is the observation of entanglement sudden death for every subsystem. The effects of various interactions such as Ising interaction, single photon exchange interaction and dipole-dipole interaction on entanglement dynamics are studied. The effects of detuning, Kerr-nonlinearity on the entanglement dynamics are investigated for every subsystem. It is noticed that proper choice of the interaction parameters associated with photon exchange interaction, dipole-dipole interaction, detuning and Kerr-nonlinearity effectively remove entanglement deaths from the dynamics.
@@ 1 Introduction
It is well known that the Jaynes-Cummings (J-C) model is one of the most successful models to describe the atom-field interaction and to study the dynamics of various quantum optical aspects. Entanglement [1] is one of the important quantities which can be investigated using the Jaynes-Cummings model. In recent years, entanglement has become ubiquitous and necessary for quantum applications. It has become a resource for quantum teleportation [2], super dense coding [3], entanglement swapping, cryptography etc. In many quantum optical systems such as in ion traps [4, 5, 6], cavity quantum electrodynamics (CQED) [7, 8, 9, 10, 11, 12, 13, 14], circuit quantum electrodynamics [15] etc., the entanglement is studied. All these physical platforms involve nontrivial interactions between atoms and fields. Jaynes-Cummings model provides a platform to study entanglements in these systems.
One of the successful extensions of the usual Jaynes-Cummings model is the double Jaynes-Cummings model (DJCM), which was first introduced by Yonac et. al., [16, 17, 18, 19, 20, 21, 22]. This model can be used to study the properties of entanglement in a more extended quantum optical systems which are mentioned before. A striking feature of the DJCM is the sudden disappearance or death of entanglement for a certain duration. This phenomenon of disappearance of entanglement is termed as entanglement sudden death (ESD). After Yonac’s work, there are studies on various systems with this model [23, 24, 25, 26]. In particular in Ref. [25], the authors have studied entanglement dynamics when atoms interact with fields in coherent and squeezed states. They have shown that the factors such as atomic spontaneous decay rate, cavity decay rate and detuning have significant effects on the entanglement sudden death. To study the effects of interaction, the authors in Ref. [24, 27] considered cavities with Ising type and photon exchange interactions between the cavities. The notable features of entanglement decay, sudden rebirth and sudden death have been investigated in the presence of intrinsic decoherence in Ref. [28]. In another work, Laha [29] has considered a double Jaynes-Cummings model to investigate the role of a beam-splitter, dipole-dipole interaction as well as Ising type interactions. The photonic modes in this work are considered to be either the photonic vacuum state, or the coherent states or the thermal states and the entanglements investigated are the qubit-qubit and the oscillator-oscillator type.
@@
After Yonac’s work on double Jaynes-Cummings model and the entanglement sudden death, various systems with double Jaynes-Cummings model have been studied; however, there are very few studies on the intensity-dependent double Jaynes-Cummings model (IDDJCM). In [39] Xie Qin and Fang Mao-Fa have investigated the entanglement dynamics of IDDJCM with two different initial radiation fields. One is a coherent state and another is a squeezed vacuum state. Motivated by these works, in this paper, we study the entanglement dynamics for both double Jaynes-Cummings and intensity-dependent double Jaynes-Cummings models with the radiation fields in squeezed coherent thermal states (SCTS), and, with the atoms in a Bell state.
The intensity-dependent double Jaynes–Cummings model provides a more realistic framework by incorporating nonlinear nature of atom–field coupling, reflecting practical scenarios such as those encountered in nonlinear optics, trapped ions, or micromaser dynamics. On the other hand, the use of squeezed coherent thermal states (SCTS) as the initial radiation field offers a richer and more physically relevant model, capturing both thermal noise, coherence, and nonclassical features like squeezing. This combination of nonlinear coupling and realistic field states opens up a deeper exploration of how this structured light interacts with matter and influences entanglement evolution. The motivation behind this study is to understand how these modifications affect entanglement robustness, entanglement sudden death phenomena, and the potential for entanglement engineering in practical quantum optical systems.
One of the main objectives of this work is to investigate the effects of squeezed photons and thermal photons on the entanglement in a coherent background for DJCM and IDDJCM. Studying the effects of different kinds of interactions on the entanglement dynamics in the system is also an objective. In this paper, the effects of diverse interactions such as single photon exchange interaction, Ising interaction and dipole-dipole interaction on entanglement are investigated. We also investigate the effects of detuning, Kerr-nonlinearity [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51] etc. Earlier, in Ref.[52], the authors have studied the effects of the squeezed and thermal photons on entanglements in Jaynes-Cummings model. Here, it was shown how the tussling between “classical noise” (thermal photons) and “quantum noise” (squeezed photons) affects the quantum optical properties of the radiation field, atomic inversion and entanglement dynamics for Jaynes-Cummings model. The work presented here is a natural extension of our previous work to investigate the effects of the perpetual tussle between thermal and squeezed photons on the entanglement dynamics of various subsystems in DJCM and IDDJCM.
@@ 2.1 Photonic state
$|,\hat{\rho}_{\text{th}}=\frac{1}{1+\bar{n}_{th}}\sum_{n=0}^{\infty}\left(\frac{\bar{n}_{th}}{\bar{n}_{th}+1}\right)^{n}\ket{n}\bra{n}$, (4)
@@
$\times\big{|}H_{q}((2Y)^{-1/2}\tilde{Z})\big{|}^{2},$ (7)
where $\pi Q(0)=R(0,0)$; $R$ is Glauber’s $R$-function[56]; (see Appendix A for more detail).
@@ 2.2 Atomic state
The double Jaynes-Cummings model involves two two-level atoms: atom A and atom B, each one interacting with a cavity mode. Atom A is placed inside cavity a and atom B is placed inside cavity b. These two-level systems are initially prepared in either as a pure state or as a mixed state. For the pure state a maximally entangled Bell state is considered. In this paper, we use the entangled atomic state of the form
@@
[…]
@@ 2.3 The Hamiltonian of the system
Density operator for the atomic state is
$|,\hat{\rho}_{\textsubscript{AB}}(0)=\ket{\psi_{\textsubscript{AB}}}\bra{\psi_{\textsubscript{AB}}}$, (13)
and initial density operator representing the field states is $\hat{\rho}_{\text{F}}(0)$. So, the density operator for the whole system can be written as
$\hat{\rho}_{\text{tot}}(0)=$ $\hat{\rho}_{\textsubscript{AB}}(0)\otimes\hat{\rho}_{\text{F}}(0).$ (14)
$=$ $\hat{\rho}_{\textsubscript{AB}}(0)\otimes\hat{\rho}^{\text{a}}_{\text{F}}(0)\otimes\hat{\rho}^{\text{b}}_{\text{F}}(0),$ (15)
where $\hat{\rho}^{\text{a}}_{\text{F}}(0)$ and $\hat{\rho}^{\text{b}}_{\text{F}}(0)$ are the density operator for the radiation field in cavities a and b respectively. Since the fields in the cavities are not entangled initially, we can express $\hat{\rho}_{\text{F}}(0)$ as $\hat{\rho}^{\text{a}}_{\text{F}}(0)\otimes\hat{\rho}^{\text{b}}_{\text{F}}(0)$.
2.4 Time evolution of the system
To get the time evolve state of the system, we will use the interaction picture. The Hamiltonian in the interaction picture with the resonant condition (i.e., $\nu-\omega=0$) can be written as
$\hat{H}_{I}=\lambda\left(\sqrt{\hat{N}_{\text{a}}}\,\hat{a}^{\dagger}\sigma_{-}^{\text{A}}+\hat{a}\sqrt{\hat{N}_{\text{a}}}\,\hat{\sigma}_{+}^{\text{A}}\right)+\lambda\left(\sqrt{\hat{N}_{\text{b}}}\,\hat{b}^{\dagger}\hat{\sigma}_{-}^{\text{B}}+\hat{b}\sqrt{\hat{N}_{\text{b}}}\,\hat{\sigma}_{+}^{\text{B}}\right),$ (16)
and similarly, it can be written for the Hamiltonian in Eq. (12).
The time evolution operator for the whole system is given by
$\hat{U}(t)=e^{\frac{-i\hat{H}_{I}t}{\hbar}},$ (17)
where $\hat{H}_{I}$ is the interaction Hamiltonian for the whole system. If there is no cross-interaction present in the system, this $\hat{U}(t)$ can be written as
$\hat{U}(t)=\hat{U}^{(1)}(t)\otimes\hat{U}^{(2)}(t),$ (18)
where
$\hat{U}^{(1)}(t)=\exp\left[\frac{-i\lambda\left(\sqrt{\hat{N}_{\text{a}}}\,\hat{a}^{\dagger}\sigma_{-}^{\text{A}}+\hat{a}\sqrt{\hat{N}_{\text{a}}}\,\hat{\sigma}_{+}^{\text{A}}\right)t}{\hbar}\right],$ (19)
@@ 2.4 Time evolution of the system
$\hat{U}^{(2)}(t)=\exp\left[\frac{-i\lambda\left(\sqrt{\hat{N}_{\text{b}}}\,\hat{b}^{\dagger}\hat{\sigma}_{-}^{\text{B}}+\hat{b}\sqrt{\hat{N}_{\text{b}}}\,\hat{\sigma}_{+}^{\text{B}}\right)t}{\hbar}\right].$ (20)
The initial state of the whole system $\hat{\rho}_{\text{tot}}(0)$ in Eq. 15 can be written as
$\hat{\rho}_{\text{tot}}(0)=\sum_{i,j}c_{ij}\,|i\rangle\langle j|\otimes\hat{\rho}^{\text{a}}_{\text{F}}(0)\otimes\hat{\rho}^{\text{b}}_{\text{F}}(0),$ (21)
where $c_{ij}$ are the coefficients containing different combinations of $\sin\theta,\cos\theta$. So, the time-evolved density operator for the whole system can be written as
$\hat{\rho}_{\text{tot}}(t)=\hat{U}(t)\left(\sum_{i,j}c_{ij}\,|i\rangle\langle j|\otimes\hat{\rho}^{\text{a}}_{\text{F}}(0)\otimes\hat{\rho}^{\text{b}}_{\text{F}}(0)\right)\hat{U}^{\dagger}(t).$ (22)
This density operator can be expanded further in atomic basis. The details are shown in Appendix B.
Computing entanglement for different subsystems:
Although, an analytical expression for the density operator corresponding to a squeezed coherent thermal state (SCTS) is highly involved and no closed-form expansion in the Fock basis is known, the entanglement dynamics of the system can still be reliably and robustly studied through numerical methods. In particular, by evolving the total density matrix and performing partial traces over the appropriate subsystems, one can investigate the entanglement between atoms, fields, and hybrid subsystems. For instance, to compute the atom-atom entanglement, we can trace out the field-field subsystems from the time evolved density operator in Eq. 22. The reduced atom-atom density operator can be written as:
$\hat{\rho}_{\textsubscript{AB}}(t)=$ $\text{Tr}_{\text{a,b}}[\hat{\rho}_{\text{tot}}(t)]$ (23)
$=$ $\sum_{i,j}\text{Tr}_{\textsubscript{a,b}}\left[\hat{U}(t)\left(|i\rangle\langle j|\otimes\hat{\rho}^{\text{a}}_{\text{F}}(0)\otimes\hat{\rho}^{\text{b}}_{\text{F}}(0)\right)\hat{U}^{\dagger}(t)\right].$ (24)
To compute the entanglement between atom A and field a, we trace out the atom B and field b degrees of freedom;
$\hat{\rho}_{\textsubscript{A,a}}(t)=\mathrm{Tr}_{\textsubscript{B,b}}[\hat{\rho}_{\text{tot}}(t)],$ (25)
and similarly for other subsystems.
@@ 2.5 Entanglement measures
where $\Lambda_{i}(i=1,2,3,4)$ are the decreasingly ordered square roots of the eigenvalues of the matrix $\hat{\rho}\left(\hat{\sigma}_{y}^{\text{A}}\otimes\hat{\sigma}_{y}^{\text{B}}\right)\hat{\rho}^{*}\\ \left(\hat{\sigma}_{y}^{\text{A}}\otimes\hat{\sigma}_{y}^{\text{B}}\right)$ and $\hat{\rho}$ is the two qubit atom-atom reduced density matrix. The value of concurrence lies in the range $0\leq C\leq 1$, where $C=0$ implies a separable state and $C=1$ denotes a maximally entangled state. Though concurrence can be used to compute entanglement in both pure and mixed states, if the system is $2\otimes 2$ dimensions. So, for higher dimensional systems we need to use other measures. In particular, when we consider the atom-field subsystem we are looking at a bipartite object, whose Hilbert space is composed of a 2-dimensional Hilbert space (associated with the atom) in a tensor product with a $\infty$-dimensional Hilbert space (associated with a continuous variable system i.e. the field. For these systems it is convenient to use the negativity [58] which is defined as
@@
3 Entanglement Dynamics in IDDJCM and DJCM with SCTS
In this section, we investigate the entanglement dynamics of an atom-field system interacting with a squeezed coherent thermal state (SCTS) of the radiation field, where the atoms are initially prepared in a Bell state. Figure 1 illustrates the entanglement dynamics for the intensity-dependent double Jaynes–Cummings model (IDDJCM). The parameters chosen for this study are $\bar{n}_{c}=2$, $\bar{n}_{th}=1.0$, and $\bar{n}_{s}=0.1,\ 0.3,\ 0.5,\ 1.0$. Previous studies on the intensity-dependent Jaynes–Cummings model have established that the system dynamics exhibit a periodic structure [39].
The chosen values $\bar{n}_{c}=2$, $\bar{n}_{\text{th}}=1$, and $\bar{n}_{s}=0.1\text{--}1.0$ ensure a physically meaningful balance between classicality and nonclassicality, while remaining experimentally realistic and numerically stable. This allows us to explore how thermal photons, coherent photons, and squeezing individually and collectively influence the entanglement dynamics in the IDDJCM framework.
From Fig. 1, it is evident that the atom-atom entanglement $C(t)$ (blue curve), as well as the negativity $N(t)$ for the three hybrid and field-field subsystems, all display periodic behavior (see Appendix C for alternative figure). Specifically, the dynamics of $C(t)$ remain largely unchanged with increasing values of squeezed photons $\bar{n}_{s}$. The duration of the ESDs varies slightly with $\bar{n}_{s}$, but the peak amplitudes remain unaffected. A similar behavior is observed for the field-field subsystem (black curve in Fig. 1), where both the shape and amplitude of $N(t)$ are invariant under changes in $\bar{n}_{s}$. The explanation for this periodic behaviour of entanglement dynamics is as follows: in IDDJCM, the atom–field coupling strength becomes a function of the photon number, such as $\lambda(n)\sim\sqrt{n}$ or a saturating form, rather than remaining constant as in the standard Jaynes–Cummings model. This intensity-dependent coupling plays a crucial role in stabilizing the entanglement dynamics. Specifically, it reduces the sensitivity of the system to fluctuations in the average photon number. In the standard model, an increase in field strength leads to proportionally faster Rabi oscillations, as the atom effectively experiences a stronger driving field. However, in the IDDJCM, the nonlinear or saturating nature of the coupling “flattens” this response, keeping the atom–field interaction effectively bounded even as the photon number increases.
As a result, characteristic features such as entanglement sudden deaths (ESDs) and the amplitude and timing of entanglement peaks remain relatively unchanged across a range of average photon numbers. This effect is particularly pronounced in subsystems like atom–atom and field–field, where the dynamics are dominated by coherent rephasing, and are less affected by photon number–induced dephasing. Thus, the intensity-dependent coupling provides a form of dynamical insulation against variations in field intensity, leading to more robust and predictable entanglement behavior.
Entanglement corresponding to the atom A–field $a$ subsystem (green curve), the amplitude of $N(t)$ increases with $\bar{n}_{s}$, while the overall shape of the dynamics and the ESD characteristics remain unchanged. Conversely, in for the atom A–field $b$ subsystem (red curve), both the peak amplitudes and the durations of the ESDs decrease as $\bar{n}_{s}$ increases. Overall, the increase in the number of squeezed photons does not significantly impact the entanglement dynamics of the field-field subsystem, and the curves for different $\bar{n}_{s}$ values largely overlap.
Figure 2 presents the entanglement dynamics for the standard double Jaynes–Cummings model (DJCM). Unlike the IDDJCM, the entanglement in DJCM does not exhibit a periodic, pulse-like structure (see Appendix C for alternative figure). In the case of atom-atom entanglement, $C(t)$ begins at its maximum value and then undergoes a sharp decline. A pronounced ESD occurs, and its duration increases with rising $\bar{n}_{s}$. The influence of thermal photons is clearly visible in this behavior. A significant distinction arises between the intensity-dependent double Jaynes–Cummings model (IDDJCM) and the standard double Jaynes–Cummings model (DJCM) in terms of entanglement behavior. In the standard DJCM, the Rabi frequency scales as $\Omega_{n}=2\lambda\sqrt{n+1}$, leading to a broad spread of frequencies associated with different photon number states. This results in dephasing due to destructive interference, thereby suppressing entanglement revivals and producing aperiodic or damped dynamics.
In [59], smaller peaks are typically observed in the dynamics of $C(t)$, for squeezed coherent states [60, 61, 53, 54, 62, 55, 63, 64, 65, 66, 67, 68, 69, 70], and the peaks around $\lambda t=8$ are particularly prominent. However, in the presence of thermal photons, these smaller peaks vanish and the heights of the larger peaks decrease as well. This suggests that thermal photons contribute to the degradation of atom-atom entanglement in the system.